OpenStudy (anonymous):

Find the derivative of F(t) = (1-7t)^6

OpenStudy (anonymous):

U Know Hw to Find the Derivative of a Composite Function..?

OpenStudy (anonymous):

(1-7t)^6 = 6 ( 1-7t)^5 * -7 = -42 (1-7t)^5

OpenStudy (anonymous):

I got that the derivative of F(t) is 6(1-7t)^5 but it says it's -42(1-7t)^5

OpenStudy (anonymous):

Where did you get the -7 from?

OpenStudy (anonymous):

i Told u It is a Composite Function...)

OpenStudy (anonymous):

Okay thanks!

OpenStudy (anonymous):

Welcome...:)